OpenStudy (anonymous):

Super unleaded gas costs 9 cents more per gallon than regular unleaded gasoline. A service station buys 3,000 gallons of super unleaded and 2,000 gallons of regular unleaded for $8,919. What is the cost per gallon of each kind of gasoline?

OpenStudy (anonymous):

Okay this is my second attempt to answer this (this website is running slow for me some reason). So let x= cost per gallon of regular unleaded and let y= cost per gallon of super unleaded. According to your first equation, you could set up an equation like so: y=x+0.09 (since the super unleaded, y, is 9 cents more than regular, x). As for the rest of your question you can set up equation where you setx*2000 to figure out the cost for the regular unleaded and y*3000 to figure out the cost for the super unleaded. Now if you add these together you should get your total cost, $8,919. So 2000*x+3000*y=8919. Since you now have two equations to work with you can use the first one and substitute it into your second equation (so for every y you see in the second equation you can replace it with "x+0.09"). So now your second equation will look like this: 2000x+3000(x+0.09)=8919. Solve for x. Once you find your x value you can go back to the first equation to find your y value.